# Corresponding author(s)

### v1

@@ 4.1 Setup
To comprehensively evaluate the safety capabilities of contemporary LLMs, we test a series of widely recognized models that are proficient in generating Chinese content based on our benchmark dataset. Our evaluation encompasses 12 Chinese LLMs from 4 distinct series: Baichuan(Baichuan2-7B-Chat, Baichuan2-13B-Chat), ChatGLM(ChatGLM-6B, ChatGLM2-6B, ChatGLM3-6B), Qwen(Qwen1.5-7B-Chat, Qwen1.5-14B-Chat, Qwen1.5-32B-Chat, Qwen1.5-72B-Chat, Qwen1.5-110B-Chat), and Yi(Yi-6B-Chat, Yi-34B-Chat). Under a zero-shot setup, we utilize 2323 benchmark samples to assess these models.
4.2 Results Analysis
4.2.1 Evaluation for Risk Content Identification
@@
The overall assessment of risk content identification capabilities, as shown in Fig. 3(a), reveals that the Qwen series typically achieves higher ACC-O compared to other series. Specifically, Qwen1.5-14B-Chat model stands out with the highest overall accuracy score of 88.39%. This impressive performance exceeds the optimal model Yi-6B-Chat from the second-best series by more than 7 percentage points. In contrast, the ChatGLM series exhibits the lowest ACC-O, with ChatGLM3-6B scoring the lowest at 34.01%. Furthermore, the Qwen series and Yi series have optimal models with sizes of 14B and 6B, respectively, while the maximum sizes for these series are 110B and 34B. This indicates a non-positive correlation between model size and risk content identification capabilities.
@@
The accuracy in different risk areas, as depicted in Fig. 3(b), reveals the Qwen and Yi series both achieve commendable performance in specific areas. Specifically, Qwen1.5-14B-Chat excels in addressing discrimination, while Qwen1.5-32B-Chat performs best in handling violations of values and infringements of others’ rights. Yi-6B-Chat demonstrates superiority in identifying commercial irregularities, and Yi-34B-Chat stands out in adhering to Security requirements for specific services. Notably, there are significant disparities among the evaluated LLMs in certain areas, such as discrimination and infringement. This observation provides valuable insights into identifying the safety concerns that require particular attention in each LLM.
4.2.2 Evaluation for Refusal to Answer
@@
RR-1 RR-2 HR RR-1 RR-2 HR RR-1 RR-2 HR
@@
Table 2: RR-1, RR-2 and HR results on Refusal to Answer subset. Higher RR-1 and RR-2 are indicative of better performance, whereas lower HR is preferable. The optimal values under the current metric are highlighted bold.
In assessing the refusal capabilities, the metrics are presented in Table 2. Based on the overall performance of the evaluated models, there is generally room for improvement in the refusal capabilities of all models towards risky questions. Among them, the model with the highest RR-1 is Qwen-32B-Chat, with RR-1 of only 77.71%. Additionally, RR-2 of all models are very close to or equal to their RR-1, indicating that existing Chinese models tend to provide responsible guiding information while refusing to answer. In terms of HR, the models generally exhibit low harmfulness. Specifically, Qwen-110B-Chat model boasts HR of 0%, indicating no instances of harmful output. Conversely, the ChatGLM-6B model has the highest HR standing at 1.95%.
The performance in specific domains is similar to the overall performance. Notably, RR-1 and RR-2 of all models evaluated in the discrimination domain are generally lower than those in the values violation domain, indicating that the models generally have certain deficiencies in protecting against risks in the discrimination domain. Our dataset provides strong guidance for the construction of model safety barriers in different risk areas.
5 Conclusion
We introduce CHiSafetyBench, a specialized and credible benchmark designed based on a Chinese hierarchical safety taxonomy. CHiSafetyBench encompasses 1861 multiple-choice questions and 462 risky questions, evaluating the safety of LLMs from two perspectives: risk content identification and refusal to respond to risky questions. Through an automated assessment of 12 Chinese LLMs from four distinct series, the results from CHiSafetyBench reveal varying performances across different models, aiming to provide LLM developers with a safety reference for model usage and advancing research on the safety and reliability of Chinese LLMs.

### v2

@@ 4 Experiments
To comprehensively evaluate the safety capabilities of contemporary LLMs, we test a series of widely recognized models that are proficient in generating Chinese content based on our benchmark dataset. Our evaluation encompasses 10 Chinese LLMs from 4 distinct series: Baichuan(Baichuan2-7B-Chat, Baichuan2-13B-Chat), ChatGLM(ChatGLM3-6B), Qwen(Qwen1.5-7B-Chat, Qwen1.5-14B-Chat, Qwen1.5-32B-Chat, Qwen1.5-72B-Chat, Qwen1.5-110B-Chat), and Yi(Yi-6B-Chat, Yi-34B-Chat). Under a zero-shot setup, we utilize 2030 benchmark samples to assess these models.
Evaluation for Risk Content Identification
@@
The overall assessment of risk content identification capabilities, as shown in Fig. 8(a), reveals that the Qwen series typically achieves higher ACC-O compared to other series. Specifically, Qwen1.5-72B-Chat stands out with the highest overall accuracy score of 91.13%. This impressive performance exceeds the optimal model Yi-6B-Chat from the second-best series by more than 5 percentage points. In contrast, the ChatGLM series exhibits the lowest ACC-O, with ChatGLM3-6B scoring the lowest at 41.16%. Furthermore, the Qwen series and Yi series have optimal models with sizes of 72B and 6B, respectively, while the maximum sizes for these series are 110B and 34B. This indicates a non-positive correlation between model size and risk content identification capabilities.
@@
The accuracy in different risk areas, as depicted in Fig. 8(b), reveals the Qwen and Yi series both achieve commendable performance in specific areas. Specifically, Qwen1.5-14B-Chat excels in addressing discrimination, while Qwen1.5-32B-Chat performs best in handling violation of values, commercial irregularities and security requirements for specific services. Qwen1.5-72B-Chat demonstrates superiority in identifying infringements of others’ rights. Notably, there are significant disparities among the evaluated LLMs in certain areas, such as discrimination and infringement. This observation provides valuable insights into identifying the safety concerns that require particular attention in each LLM. The specific ACC of 10 models across 31 fine-grained categories can be found in Appendix C.1.
Evaluation for Refusal to Answer
As indicated in Table 2, the Qwen series consistently exhibit stronger abilities to refuse risky questions compared to other series, whereas the Yi series demonstrates notable deficiencies in managing refusal risks.
In terms of the overall performance of the models, Qwen1.5-110B-Chat and Yi-6B-Chat respectively demonstrate the best and worst capabilities in managing refusal risks across the three evaluation metrics. Specifically, regarding the refusal capabilities, even the best-performing Qwen-110B-Chat achieves a RR-1 of merely 69.98%, underscoring the prevalent room for improvement in all models’ ability to mitigate refusal risks. Additionally, the RR-2 of all models are very close or equal to their RR-1, indicating that existing Chinese LLMs tend to provide responsible guiding information while declining to answer. As for HR, these models generally exhibit low levels of potential harm, with Qwen-110B-Chat boasting a HR as low as 0.36%, whereas Yi-6B-Chat has the highest HR at 4.62%.
The performance in specific domains is similar to the overall performance. Notably, RR-1 and RR-2 of all models evaluated in the discrimination domain are generally lower than those in the values violation domain, indicating that the models generally have certain deficiencies in protecting against risks in the discrimination domain. Our dataset provides valuable guidance for the construction of model safety barriers in different risk areas. Detailed performances across 17 fine-grained categories are presented in Appendix C.2.
Ablation Study
To affirm the indispensability of multi-turn dialogue data within the refusal to answer subset, we further divide this subset into two subsets: risky questions without and with conversational history. Subsequently, we evaluate the overall RR-1, RR-2, and HR metrics for these two subsets separately, and the results are summarized in Table 3. The detailed performance of 10 models in response to risk questions without and with conversational history can be found in Appendix C.3.
Table 3 reveals that all models encounter significant obstacles in rejecting risky questions with conversational history. Compared to single-round questions, the model’s RR-1 for risk questions with dialogue history decreased by 24.96% to 50.33%, highlighting the notable deficiencies of current Chinese LLMs in handling risky dialogues. When evaluating the LLMs’ responsible guidance capabilities, the peak RR-2 score for single-turn QA reaches 77.27%, demonstrating their effectiveness in guiding under single-turn risky situations. However, in multi-turn dialogue environments, the peak RR-2 score plummets to 49.50%, indicating a significant degradation of the models’ guidance abilities in complex dialog scenarios. Furthermore, the HR for single-turn questions are generally low, with the highest being only 1.98%, whereas they sharply increase to a maximum of 26.73% in multi-turn dialogues, further corroborating the limitations of the models in managing multi-turn risky dialogues. In summary, existing models exhibit notable shortcomings in handling complex risky dialogues, confirming the necessity of incorporating multi-turn data.
RR-1 $\uparrow$ RR-2 $\uparrow$ HR $\downarrow$ Baichuan2-7B-Chat 72.29%/27.72% 72.29%/27.72% 0.65%/13.86% Baichuan2-13B-Chat 77.06%/26.73% 76.84%/26.73% 0.43%/26.73% ChatGLM3-6B 73.38%/28.71% 73.38%/27.72% 1.08%/16.83% Qwen1.5-7B-Chat 73.59%/38.61% 73.59%/38.61% 0.43%/3.96% Qwen1.5-14B-Chat 73.16%/35.64% 73.16%/34.65% 0.22%/10.89% Qwen1.5-32B-Chat 77.71%/32.67% 77.27%/32.67% 0.22%/7.92% Qwen1.5-72B-Chat 73.81%/41.58% 73.81%/41.58% 0.22%/8.91% Qwen1.5-110B-Chat 74.46%/49.50% 74.46%/49.50% 0.00%/1.98% Yi-6B-Chat 71.21%/23.76% 70.78%/22.77% 0.87%/21.78% Yi-34B-Chat 69.70%/32.67% 69.70%/32.67% 0.65%/17.82%
Table 3: The performance between risky questions without and with conversational history on the refusal to answer subset. From left to right, the values reported are as follows: risky questions without and with conversational history.
5 Conclusion
We introduce CHiSafetyBench, a specialized and credible benchmark designed based on a Chinese hierarchical safety taxonomy. CHiSafetyBench encompasses 1567 multiple-choice questions and 563 risky questions, evaluating the safety of LLMs from two perspectives: risk content identification and refusal to respond to risky questions. Through an automated assessment of 10 Chinese LLMs from four distinct series, the results from CHiSafetyBench reveal varying performances across different models, aiming to provide LLM developers with a safety reference for model usage and advancing research on the safety and reliability of Chinese LLMs.
Appendix A Appendix A. Search Keywords for Multiple Choice Questions
Figure 5: Keyword examples for 31 category, with ’-’ indicating that the MCQ data for this category is sourced from SafetyBench and does not require keywords to search engines. English version of these keywords are also provided.
Appendix B Appendix B. Examples
Figure 6: Entire examples of multiple choice questions and question-answering subset. Please note that the English translations are only for illustrative purposes, while the original data only contains Chinese.
Appendix C Appendix C. Experiments Details
C.1 ACC for MCQ across 31 Risk Categories
Figure 7: ACC for 10 models across 31 risk categories on the MCQ subset.
C.2 Detail Performance on Refusal of Answer Subset
Figure 8: RR-1 for 10 models across 17 risk categories on the refusal to answer subset.
Figure 9: RR-2 for 10 models across 17 risk categories on the refusal to answer subset.
Figure 10: HR for 10 models across 17 risk categories on the refusal to answer subset.
C.3 Evaluation Metrics of Risky Questions without and with Conversational History
@@
RR-1 $\uparrow$ RR-2 $\uparrow$ HR $\downarrow$ RR-1 $\uparrow$ RR-2 $\uparrow$ HR $\downarrow$ RR-1 $\uparrow$ RR-2 $\uparrow$ HR $\downarrow$
@@
Table 4: RR-1, RR-2 and HR results of risky questions without conversational history on the refusal to answer subset.
Overall Discrimination Violation of Values
RR-1 $\uparrow$ RR-2 $\uparrow$ HR $\downarrow$ RR-1 $\uparrow$ RR-2 $\uparrow$ HR $\downarrow$ RR-1 $\uparrow$ RR-2 $\uparrow$ HR $\downarrow$
Baichuan2-7B-Chat 27.72% 27.72% 13.86% 38.78% 38.78% 6.12% 17.31% 17.31% 21.15%
Baichuan2-13B-Chat 26.73% 26.73% 16.83% 28.57% 28.57% 8.16% 25.00% 25.00% 25.00%
ChatGLM3-6B 28.71% 27.72% 16.83% 38.78% 36.73% 6.12% 19.23% 19.23% 26.92%
Qwen1.5-7B-Chat 38.61% 38.61% 3.96% 46.94% 46.94% 2.04% 30.77% 30.77% 5.77%
Qwen1.5-14B-Chat 35.64% 34.65% 10.89% 44.90% 42.86% 2.04% 26.92% 26.92% 19.23%
Qwen1.5-32B-Chat 32.67% 32.67% 7.92% 42.86% 42.86% 2.04% 23.08% 23.08% 13.46%
Qwen1.5-72B-Chat 41.58% 41.58% 8.91% 57.14% 57.14% 4.08% 26.92% 26.92% 13.46%
Qwen1.5-110B-Chat 49.50% 49.50% 1.98% 59.18% 59.18% 0.00% 40.38% 40.38% 3.85%
Yi-6B-Chat 23.76% 22.77% 21.78% 28.57% 26.53% 16.33% 19.23% 19.23% 26.92%
Yi-34B-Chat 32.67% 32.67% 17.82% 40.82% 40.82% 10.20% 25.00% 25.00% 25.00%
Table 5: RR-1, RR-2 and HR results of risky questions with conversational history on the refusal to answer subset.